# Active Fine-Tuning of Generalist Policies

### v1

@@ 6 Discussion
Naturally, active multi-task fine-tuning has several limitations. When coupled with neural networks, the algorithm relies on uncertainty estimation techniques, which remain an open problem. While the approximation we leverage is informative in our experiments, AMF could benefit if large pre-trained policies would allow other off-the-shelf uncertainty quantification techniques (e.g., through model ensembling during pre-training). Second, we found the performance of AMF to depend naturally on the pre-training data distribution. While AMF induces efficient learning for skewed pre-training distributions, it naturally brings more modest gains when the pre-trained policy is equally capable for all tasks, and uniform task sampling is sufficient.
On top of addressing the current limitations, this work suggests multiple interesting directions. An extensive empirical evaluation of active fine-tuning with large-scale generalist policies is clearly desirable, but remains infeasible at the moment due to the scarce availability of open-source models and benchmarks. Another future research direction would involve direct estimation of the RL objective, thus removing the dependence on non-equivalent BC proxy objectives.
@@ Acknowledgments
Reproducibility Statement
Appendix I describes implementation details and hyperparameters across all experiments. We additionally open-source a clean and commented implementation of AMF-GP and AMF-NN at the anonymous website.
@@ A.2 Proof
We first prove that, under Assumptions 2 and 3, the effect of stochasticity on the mutual information at step $n$ is bounded.
Lemma 1.
Let Assumptions 2 and 3 hold. Fix a sequence of tasks $c_{1:n-1}$ and consider two arbitrary sequences of trajectories $\tau_{1:n-1}$ and $\tau^{\prime}_{1:n-1}$ sampled from $\bm{\tau}(c_{1:n-1})$. Fix one state-task pair $(s,c)\in\mathcal{S\times C}$, one task $c_{n}\in\mathcal{C}$ and one trajectory $\tilde{\tau}\sim\bm{\tau}(c_{n})$. Let $\epsilon_{n}=8H^{\frac{3}{2}}(1+\max(L_{P},L_{\pi}))^{H}\max(\epsilon_{0},\epsilon_{\pi},\epsilon_{P})$. The difference in mutual information when conditioning on the two sequences of trajectories can be bounded:
@@
Under Assumption 3 it is sufficient to show that stochasticity in the MDP does not cause the demonstrator’s trajectories to deviate excessively. This is a direct consequence of smoothness and boundedness, which we assume in Assumption 2, and can be shown by induction. Let us fix a task $c_{n}\in\mathcal{C}$ and consider two trajectories $\tau,\tau^{\prime}\sim\bm{\tau}(c_{n})$. For the two initial states $(s_{0},s^{\prime}_{0})$, boundedness of the initial state distribution $\mu_{0}$ implies that $\|s_{0}-s^{\prime}_{0}\|_{2}\leq\epsilon_{\mu_{0}}$. Now, assuming that the distance between two states $(s_{t},s^{\prime}_{t})$ is bounded as $\|s_{t}-s^{\prime}_{t}\|_{2}\leq\epsilon_{t}$, we have that
@@
where Lemma 9 was used in (i) and (ii); Assumption 2 and the fact that $c_{t}=c^{\prime}_{t}$ were used through the rest of the derivation. The recurrence relation can be easily unrolled as
@@
thus bounding the L2 distances between states at each step of the trajectory $\epsilon_{t}=\|s_{t}-s^{\prime}_{t}\|_{2}$. We note that the distance between actions can also be easily bound by Lemma 9: $\|a_{t}-a^{\prime}_{t}\|_{2}\leq L_{\pi}\epsilon_{t}+2\epsilon_{\pi}$. This can in turn be related to distances over trajectories. Let us fix $c_{1:n-1}\in\mathcal{C}$ and consider $\tau_{1:n-1},\tau^{\prime}_{1:n-1}\sim\bm{\tau}(c_{1:n-1})$. We have that
@@
Having obtained an upper bound on the distance between sequences of trajectories, the result follows naturally from smoothness of mutual information according to Assumption 3.
@@
This can be leveraged to show that the expected mutual information is sublinear in the number of rounds $n$. From this point, we overload the notation and allow policies (e.g. $\bm{\pi}$) to map vector to random vectors, that is $\bm{\pi}((x_{0},\dots,x_{n-1}))=(\bm{\pi}(x_{0}),\dots,\bm{\pi}(x_{n-1}))$ for $(x_{0},\dots,x_{n-1})\in(\mathcal{S\times C})^{n}$.
Lemma 2.
Under Assumptions 1 and 4, if $(c_{0},\dots,c_{n})$ follows the criterion in Equation 6, then $\Gamma_{n}\leq\frac{H}{n}\gamma_{(Hn)}$, where $\gamma_{(Hn)}=\max_{\begin{subarray}{c}X\subseteq\mathcal{S}\times\mathcal{C}\\ |X|\leq Hn\end{subarray}}\mathcal{I}(\bm{\pi}(\mathcal{S\times C});\bm{\tilde{\pi}}(X))$.
@@
where (i) follows from Assumption 4, (ii) follows from Equation 6, (iii) is due to the chain rule of mutual information. We note that $\gamma_{n}=\max_{X\subseteq\mathcal{S}\times\mathcal{C},\;|X|\leq n}\mathcal{I}(\bm{\pi}(\mathcal{S\times C});\bm{\tilde{\pi}}(X))$ is sublinear for a large class of GPs. In this cases, a looser upper bound would be $H^{2}\frac{\gamma_{n}}{n}$. ∎
@@
Lemma 3.
(Uniform convergence of marginal variance, following Hübotter et al. (2024)) Under Assumption 1, 2 and 3, for any $n\geq 0$ and $(s,c)\in\mathcal{S\times C}$,
@@
where (i) follows from Lemma 8 and monotonicity of variance, (ii) holds as the state $s$ is within the support of $\bm{\tau}(\cdot|c)$, and (iii) follows from Lemma 1 as the difference between the expected mutual information and the mutual information for a realized trajectory is less than the difference in mutual information for two arbitrary realized trajectories. ∎
@@
We note that $\beta_{n}(\delta)$ depends on $\gamma_{(Hn)}$ as $Hn$ samples from the demonstrator’s policy are collected up to round $n$. Combining Lemmas 3 and 4 we easily get for all $(s,c)\in\mathcal{S\times C}$ and $n\geq 0$ with probability $1-\delta$:
@@
Theorem 3.
(Performance guarantees for active multi-task BC) Let Assumptions 2, 3 and 4 hold. Pick $\delta\in(0,1)$. Assume that $\pi^{\star}$ lies in the RKHS $\mathcal{H}_{k}(\mathcal{C})$ of the kernel $k$ with norm $||\pi^{\star}||_{k}<\infty$, the noise $\epsilon_{n}$ is conditionally $\rho$-sub-Gaussian, and $\gamma_{n}$ is sublinear in $n$. If each demonstrated task is selected according to the criterion in Equation 1, then with probability at least $1-\delta$ the performance difference between the expert policy $\pi^{\star}$ and the imitator policy $\pi_{n}$ after $n$ demonstrations can be upper bounded:
$J^{\pi^{\star}}-J^{\pi_{n}}\leq\frac{\sqrt{2}L_{Q^{\pi^{\star}}}\|\bar{\sigma}\|_{1}}{\tau_{\text{min}}(1-\gamma)}\Big{(}(1+\epsilon_{n})\beta_{n}^{{}^{\prime}2}(\delta)\Gamma_{n}\Big{)}^{\frac{1}{2}}=O(\gamma_{(Hn)}n^{-\frac{1}{2}}),$
@@
where (i) is due to the fact that $J^{\pi^{\star}}\geq J^{\pi}$ for any policy $\pi$, and the expectation fades due to uniform convergence. The only terms with a dependency on $n$ are $\beta^{\prime}_{n}(\delta)=O(\gamma_{(Hn)}^{\frac{1}{2}})$ and $\Gamma_{n}=O(\gamma_{(Hn)}n^{-1})$, which can be combined in the asymptotic notation in the Theorem. If $\gamma_{n}=O(\log n)$, then $J^{\pi^{\star}}-J^{\pi}=O(n^{-\frac{1}{2}}\log n)\stackrel{{\scriptstyle n\to\infty}}{{\to}}0$. For a summary of magnitudes of $\gamma_{n}$ for common kernels, we refer to Table 3 in Hübotter et al. (2024). ∎
@@ Appendix B Guarantees in non-Lipschitz MDPs
The main result reported in Theorem 3 provides anytime guarantees on the agent’s performance, assuming smoothness in the MDP. However, it is possible to replace this assumption with a weaker one, at the cost of only retaining guarantees in expectation. This weaker version of the theorem can be retrieved by simply assuming smoothness on the noise, rather than on the MDP, and leveraging results recently presented by Maran et al. (2023).
Assumption 5.
@@
This assumption is satisfied by a large class of Gaussian and sub-Gaussian distributions (Maran et al., 2023). We can build upon Assumption 4 and Lemma 2, and start by providing a weaker version of Lemma 3.
Lemma 6.
(Uniform convergence of marginal variance in expectation) Under Assumption 1, for any $n\geq 0$ and $(s,c)\in\mathcal{S\times C}$,
@@
We resume from Inequality 60 in the proof of Lemma 3:
@@
Having bounded variance at each round, this time in expectation, we can invoke Lemma 4 to bound the expected distance to the optimal policy with high probability:
@@
Lemma 7.
Let $\pi^{\star}$, $\bm{\tilde{\pi}}$ and $\pi$ denote the expert, noisy expert and imitator policy, respectively. If Assumption 5 holds, then:
@@ Appendix E Additional results for AMF-NN
Results in Figure 4 are computed over two representative pre-training distributions: one allocating pre-training demonstrations uniformly over all tasks, the other one only demonstrating the first two tasks. We report these results again, and compare them with those for several other pre-training distributions. In particular, we evaluate a family of skewed priors which are only trained on one or two tasks. Results are reported for FrankaKitchen in Figure 9 and for Metaworld in Figure 10, and are consistent with patterns observed in Figure 4.
0510Rounds0.40.50.60.7Multi-task success rateKitchen, uniform
@@
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
@@
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
@@
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
@@
0510Rounds0.20.40.60.8Multi-task success rateKitchen, skewed
@@
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[16,0,0,0,0]$
  AMF-NN ($\sigma=1e^{-3}$)    Uniform    Uniform, no parameter isolation
@@
0510Rounds0.40.50.60.70.8Multi-task success rateMetaworld, uniform
@@
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
@@
0510Rounds0.20.40.60.8Multi-task success rateMetaworld, skewed
@@
0510Rounds0.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
@@
0510Rounds0.20.40.60.8Multi-task success rateMetaworld, skewed
$[16,0,0,0]$
  AMF-NN ($\sigma=1e^{-3}$)   Uniform    Uniform, no parameter isolation
@@
Moreover, we also report an ablation over the choice of batch size for Metaworld in Figure 11, thus complementing the one reported in Section 5.3. We observe a slight upward trend favoring larger batch size in the case of an uniform prior, but smaller batches remain overall desirable.
34567Area under success rate curveMetaworld, uniform
34567Area under success rate curveMetaworld, skewed
$B\!=1$
$B\!=2$
$B\!=4$
@@
Uniform, no
parameter isolation
Figure 11: Influence of batch size $B$ over area under success rate curve with a budget of 20 demonstrations in Metaworld.
Appendix F Uncertainty ablation
Section 5.4 evaluates alternative uncertainty quantification schemes in FrankaKitchen for two pre-training distributions. This Section extends these results to include results for Metaworld, and for several other pre-training settings (see Figures 12 and 13). Results are consistent with those so far reported, suggesting that loss gradient embeddings are an important component for the empirical performance of AMF-NN.
0510Rounds0.550.600.650.700.75Multi-task success rateKitchen, uniform
$[3,3,3,3,3]$
0510Rounds0.40.50.60.70.8Multi-task success rateKitchen, skewed
$[8,8,0,0,0]$
0510Rounds0.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[0,8,8,0,0]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[0,0,8,8,0]$
0510Rounds0.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[0,0,0,8,8]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
$[16,0,0,0,0]$

### v3

@@ 6 Discussion
Naturally, active multi-task fine-tuning has several limitations. When coupled with neural networks, the algorithm relies on uncertainty estimation techniques, which remain an open problem. While the approximation we leverage is informative in our experiments, AMF could benefit if large pre-trained policies would allow other off-the-shelf uncertainty quantification techniques (e.g., through model ensembling during pre-training). Second, we found the performance of AMF to depend naturally on the pre-training data distribution. While AMF induces efficient learning for mismatched pre-training distributions, it naturally brings more modest gains when the pre-trained policy is equally capable for all tasks, and uniform task sampling is sufficient.
On top of addressing the current limitations, this work suggests multiple interesting directions. An extensive empirical evaluation of active fine-tuning with large-scale generalist policies is clearly desirable, but remains infeasible at the moment due to the scarce availability of open-source calibrated benchmarks. Another future research direction would involve direct estimation of the RL objective, thus removing the dependence on non-equivalent BC proxy objectives.
@@ Acknowledgments
Impact Statement
This paper formally introduces a framework for active data selection in MDPs, and investigates its applicability in diverse settings. Although this direction as a whole might eventually have an effect on labor and automation, we do not foresee a direct impact of our work in terms of societal consequences.
@@ A.2 Proof
We first prove that, under Assumptions A.2 and A.3, the effect of stochasticity on the mutual information at step $n$ is bounded.
Lemma A.4.
Let Assumptions A.2 and A.3 hold. Fix a sequence of tasks $c_{1:n-1}$ and consider two arbitrary sequences of trajectories $\tau_{1:n-1}$ and $\tau^{\prime}_{1:n-1}$ sampled from $\bm{\tau}(c_{1:n-1})$. Fix one state-task pair $(s,c)\in\mathcal{S\times C}$, one task $c_{n}\in\mathcal{C}$ and one trajectory $\tilde{\tau}\sim\bm{\tau}(c_{n})$. Let $\epsilon_{n}=8H^{\frac{3}{2}}(1+\max(L_{P},L_{\pi}))^{H}\max(\epsilon_{0},\epsilon_{\pi},\epsilon_{P})$. The difference in mutual information when conditioning on the two sequences of trajectories can be bounded:
@@
Under Assumption A.3 it is sufficient to show that stochasticity in the MDP does not cause the demonstrator’s trajectories to deviate excessively. This is a direct consequence of smoothness and boundedness, which we assume in Assumption A.2, and can be shown by induction. Let us fix a task $c_{n}\in\mathcal{C}$ and consider two trajectories $\tau,\tau^{\prime}\sim\bm{\tau}(c_{n})$. For the two initial states $(s_{0},s^{\prime}_{0})$, boundedness of the initial state distribution $\mu_{0}$ implies that $\|s_{0}-s^{\prime}_{0}\|_{2}\leq\epsilon_{\mu_{0}}$. Now, assuming that the distance between two states $(s_{t},s^{\prime}_{t})$ is bounded as $\|s_{t}-s^{\prime}_{t}\|_{2}\leq\epsilon_{t}$, we have that
@@
where Lemma N.2 was used in (i) and (ii); Assumption A.2 and the fact that $c_{t}=c^{\prime}_{t}$ were used through the rest of the derivation. The recurrence relation can be easily unrolled as
@@
thus bounding the L2 distances between states at each step of the trajectory $\epsilon_{t}=\|s_{t}-s^{\prime}_{t}\|_{2}$. We note that the distance between actions can also be easily bound by Lemma N.2: $\|a_{t}-a^{\prime}_{t}\|_{2}\leq L_{\pi}\epsilon_{t}+2\epsilon_{\pi}$. This can in turn be related to distances over trajectories. Let us fix $c_{1:n-1}\in\mathcal{C}$ and consider $\tau_{1:n-1},\tau^{\prime}_{1:n-1}\sim\bm{\tau}(c_{1:n-1})$. We have that
@@
Having obtained an upper bound on the distance between sequences of trajectories, the result follows naturally from smoothness of mutual information according to Assumption A.3.
@@
This can be leveraged to show that the expected mutual information is sublinear in the number of rounds $n$. From this point, we overload the notation and allow policies (e.g., $\bm{\pi}$) to map vector to random vectors, that is $\bm{\pi}((x_{0},\dots,x_{n-1}))=(\bm{\pi}(x_{0}),\dots,\bm{\pi}(x_{n-1}))$ for $(x_{0},\dots,x_{n-1})\in(\mathcal{S\times C})^{n}$.
Lemma A.6.
Under Assumptions A.1 and A.5, if $(c_{0},\dots,c_{n})$ follows the criterion in Equation 6, then $\Gamma_{n}\leq\frac{H}{n}\gamma_{(Hn)}$, where $\gamma_{(Hn)}=\max_{\begin{subarray}{c}X\subseteq\mathcal{S}\times\mathcal{C}\\ |X|\leq Hn\end{subarray}}\mathcal{I}(\bm{\pi}(\mathcal{S\times C});\bm{\tilde{\pi}}(X))$.
@@
where (i) follows from Assumption A.5, (ii) follows from Equation 6, (iii) is due to the chain rule of mutual information. We note that $\gamma_{n}=\max_{X\subseteq\mathcal{S}\times\mathcal{C},\;|X|\leq n}\mathcal{I}(\bm{\pi}(\mathcal{S\times C});\bm{\tilde{\pi}}(X))$ is sublinear for a large class of GPs. In this cases, a looser upper bound would be $H^{2}\frac{\gamma_{n}}{n}$. ∎
@@
Lemma A.7.
(Uniform convergence of marginal variance, following Hübotter et al. (2024)) Under Assumption A.1, A.2 and A.3, for any $n\geq 0$ and $(s,c)\in\mathcal{S\times C}$,
@@
where (i) follows from Lemma N.1 and monotonicity of variance, (ii) holds as the state $s$ is within the support of $\bm{\tau}(\cdot|c)$, and (iii) follows from Lemma A.4 as the difference between the expected mutual information and the mutual information for a realized trajectory is less than the difference in mutual information for two arbitrary realized trajectories. ∎
@@
We note that $\beta_{n}(\delta)$ depends on $\gamma_{(Hn)}$ as $Hn$ samples from the demonstrator’s policy are collected up to round $n$. Combining Lemmas A.7 and A.8 we easily get for all $(s,c)\in\mathcal{S\times C}$ and $n\geq 0$ with probability $1-\delta$:
@@
Theorem A.10.
(Performance guarantees for active multi-task BC) Let Assumptions A.2, A.3 and A.5 hold. Pick $\delta\in(0,1)$. Assume that $\pi^{\star}$ lies in the RKHS $\mathcal{H}_{k}(\mathcal{C})$ of the kernel $k$ with norm $||\pi^{\star}||_{k}<\infty$, the noise $\epsilon_{n}$ is conditionally $\rho$-sub-Gaussian, and $\gamma_{n}$ is sublinear in $n$. If each demonstrated task is selected according to the criterion in Equation 1, then with probability at least $1-\delta$ the performance difference between the expert policy $\pi^{\star}$ and the imitator policy $\pi_{n}$ after $n$ demonstrations can be upper bounded:
$J^{\pi^{\star}}-J^{\pi_{n}}\leq\frac{\sqrt{2}L_{Q^{\pi^{\star}}}\|\bar{\sigma}\|_{1}}{\tau_{\text{min}}(1-\gamma)}\Big{(}(1+\epsilon_{n})\beta_{n}^{{}^{\prime}2}(\delta)\Gamma_{n}\Big{)}^{\frac{1}{2}}=O(\gamma_{(Hn)})/\sqrt{n},$
@@
where (i) is due to the fact that $J^{\pi^{\star}}\geq J^{\pi}$ for any policy $\pi$, and the expectation fades due to uniform convergence. The only terms with a dependency on $n$ are $\beta^{\prime}_{n}(\delta)=O(\gamma_{(Hn)}^{\frac{1}{2}})$ and $\Gamma_{n}=O(\gamma_{(Hn)})/n$, which can be combined in the asymptotic notation in the Theorem. If $\gamma_{n}=O(\log n)$, then $J^{\pi^{\star}}-J^{\pi}=O(\log n)/\sqrt{n}\stackrel{{\scriptstyle n\to\infty}}{{\to}}0$. For a summary of magnitudes of $\gamma_{n}$ for common kernels, we refer to Table 3 in Hübotter et al. (2024). ∎
@@ Appendix B Guarantees in non-Lipschitz MDPs
The main result reported in Theorem A.10 provides anytime guarantees on the agent’s performance, assuming smoothness in the MDP. However, it is possible to replace this assumption with a weaker one, at the cost of only retaining guarantees in expectation. This weaker version of the theorem can be retrieved by simply assuming smoothness on the noise, rather than on the MDP, and leveraging results recently presented by Maran et al. (2023).
Assumption B.1.
@@
This assumption is satisfied by a large class of Gaussian and sub-Gaussian distributions (Maran et al., 2023). We can build upon Assumption A.5 and Lemma A.6, and start by providing a weaker version of Lemma A.7.
Lemma B.2.
(Uniform convergence of marginal variance in expectation) Under Assumption A.1, for any $n\geq 0$ and $(s,c)\in\mathcal{S\times C}$,
@@
We resume from Inequality 60 in the proof of Lemma A.7:
@@
Having bounded variance at each round, this time in expectation, we can invoke Lemma A.8 to bound the expected distance to the optimal policy with high probability:
@@
Lemma B.3.
Let $\pi^{\star}$, $\bm{\tilde{\pi}}$ and $\pi$ denote the expert, noisy expert and imitator policy, respectively. If Assumption B.1 holds, then:
@@ Appendix E Additional results for AMF-NN
Results in Figure 4 are computed over two representative pre-training distributions: one allocating pre-training demonstrations uniformly over all tasks, the other one only demonstrating the first two tasks. We report these results again, and compare them with those for several other pre-training distributions in which tasks have been shuffled. Results are reported for FrankaKitchen in Figure 9 and for Metaworld in Figure 9, and are consistent with patterns observed in Figure 4.
01020Rounds0.20.40.60.8Multi-task success rateKitchen, mismatch
@@
01020Rounds0.20.40.60.8Multi-task success rateKitchen, mismatch
@@
01020Rounds0.20.40.60.8Multi-task success rateKitchen, mismatch
@@
01020Rounds0.30.40.50.60.70.80.9Multi-task success rateKitchen, mismatch
@@
01020Rounds0.600.650.700.750.800.85Multi-task success rateKitchen, no mismatch
@@
  AMF-NN    Uniform, with adaptive prior    Uniform
@@
01020Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, mismatch
@@
01020Rounds0.30.40.50.60.70.8Multi-task success rateMetaworld, mismatch
@@
01020Rounds0.40.50.60.7Multi-task success rateMetaworld, mismatch
@@
01020Rounds0.30.40.50.6Multi-task success rateMetaworld, no mismatch
@@
  AMF-NN ($\sigma=1e^{-3}$)   Uniform, with adaptive prior    Uniform
@@
Appendix F AMF with generalist policies
  AMF-NN
  Uniform, with adaptive prior
@@ Appendix F AMF with generalist policies
024Rounds40506070Multi-task success rateWidowX
This section investigates scaling our evaluation to recently published open-source generalist policies. For this purpose, we choose Octo (Octo Model Team et al., 2024). This model relies on a transformer backbone for integrating multimodal information (in the form of state sensors, camera images and text or RGB task descriptions), and uses a diffusion-based policy head for action prediction (Chi et al., 2023). For computational reasons, we will focus on fine-tuning the action head alone. Octo is pre-trained on a large-scale real-world robotic dataset (Collaboration, 2023), and is thus designed for inference on physical hardware. Nonetheless, a recently proposed evaluation suite enables simulated evaluations that statistically correlate with real-world results (Li et al., 2024). We thus collect rollouts from a pre-trained Octo agent on the WidowX tasks, and filter them to only include successes, akin to self-distillation schemes (Bousmalis et al., 2024). On availability of such self-supervised demonstrations, we then apply AMF-NN for $5$ iterations, providing $2$ demonstrations in each round. The results are reported in Figure 10. As all evaluation tasks are largely demonstrated in the pre-training dataset (Collaboration, 2023), and the evaluation is significantly noisier with respect to other benchmarks, we find that the performance of AMF-NN falls within confidence intervals of uniform task collection, confirming the trend we observed for uniform pre-training distributions in Figure 4. Nonetheless, we observe that it constitutes an effective method for data selection, and can be applied as a drop-in replacement for fine-tuning of off-the-shelf models. [^3] This evaluation also reports an interesting trend, that is a vast reduction in catastrophic forgetting, to the point that an adaptive prior is almost not necessary. This anecdotal evidence can be seen as an instance of a general trend of mitigated catastrophic forgetting in large models (Ramasesh et al., 2022).
Appendix G Uncertainty ablation
01020Rounds0.30.40.50.6Multi-task success rateMetaworld, no mismatch
01020Rounds0.40.50.60.70.8Multi-task success rateMetaworld, mismatch